### v1

@@ 4.2 Disentangled Causal and Confounding Features
In the previous subsection, the entity and relational representations learned based on GNN contained causal and confounding factors, and we separated them at the presentation level, which provided a solution to the previously mentioned problem of not being able to separate these two features at the data level. To do this, we introduce a decoupling module to decouple causal and confounding features. Taking the entity embedding matrix as an example, it is represented as follows:
@@
$\textbf{E}_{C}=\textbf{E}\odot\textbf{D}_{E,C},\textbf{E}_{N}=\textbf{E}\odot\textbf{D}_{E,N},$ (7)
We want the two embeddings learned from the decoupling module to be as independent as possible, which is essential to accurately separate causal and confounding features [4]. Mutual information is a basic quantity to measure the nonlinear correlation of two random variables. Minimizing mutual information is a feasible scheme to decouple causal features from confounding features. Specifically, we implement this process with contrastive log-ratio upper-bound MI estimator [7, 31], which utilizes variational distributions q and a neural network to approximate the true distribution. We define the objective function as follows:
$\mathcal{L}_{mi}=\mathbb{E}_{p(\textbf{E}_{C},\textbf{E}_{N})}\left[\log q_{\theta}(\textbf{E}_{N}|\textbf{E}_{C})\right]-\mathbb{E}_{p(\textbf{E}_{C})}\mathbb{E}_{p(\textbf{E}_{N})}\left[\log q_{\theta}(\textbf{E}_{N}|\textbf{E}_{C}^{\prime})\right].$ (8)
We conduct the same operation with relation embedding decoupling, after which we get $\textbf{R}_{C}$ and $\textbf{R}_{N}$.
@@ 4.3 Temporal Gap Guided Decoder
After the causal and confounding embeddings of entities and relations in the derived data, we use a specially crafted decoder to determine the likelihood score of potential entities and relations. Events or facts in a data stream may span different periods. For example, major political events may occur in rapid succession over a short period, while certain rare natural phenomena may occur sporadically and at longer intervals. With this in mind, it is reasonable to consider the time intervals of events to get an accurate picture of their temporal relationship. The key to the design of our decoder is the time interval vector, which guides the decoding process to consider the event time interval while calculating the fraction. Formulaic as:
$\mathbf{t}_{s}=\bm{\alpha}_{s}t+\bm{\beta}_{s},\ \mathbf{t}_{l}=\bm{\alpha}_{l}t+\bm{\beta}_{l}.$ (9)
Here, $\bm{\alpha}_{s},\bm{\beta}_{s},\bm{\alpha}_{l}$, and $\bm{\beta}_{l}$ signify learnable parameters. Adopting ConvTransE as our decoder, we introduce four variables, which traverse a one-dimensional convolutional layer followed by a fully connected layer, culminating in the extraction of a probability vector encompassing all entities. This process is mathematically articulated as:
$\mathbf{p}_{C}\left(e_{o}\mid e_{s},r,t\right)=\operatorname{ReLU}\left(\operatorname{ConvTransE\left(\mathbf{e}_{s,C,t},\mathbf{r}_{C,t},\mathbf{t}_{s},\mathbf{t}_{l}\right)}\right)\mathbf{E}_{C,t},$ (10)
@@ 4.4 Causal Intervention and Training Objective
Causal-based embedding learns the intrinsic causes that cause events to occur, so the inference results obtained from causal-based embedding are expected to yield reasonable input results. We define the supervised classification loss as follows:
@@
$\mathcal{L}_{E,N}=\frac{1}{\lvert\textbf{E}_{N,t}\rvert}\sum_{(e_{s},r,e_{o},t)\in\mathcal{G}}\operatorname{KL}\left(\textbf{y}_{u},\ \log\mathbf{p}_{N}\left(e_{o}\mid e_{s},r,t\right)\right),$ (12)
@@
$\mathbf{p}_{I}\left(e_{o}\mid e_{s},r,t\right)$ $=\operatorname{ReLU}\left(\operatorname{ConvTransE\left(\mathbf{e}_{s,I,t},\mathbf{r}_{I,t},\mathbf{t}_{s},\mathbf{t}_{l}\right)}\right)\mathbf{E}_{I,t},$ (14)
@@
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are designated as hyper-parameters, and the first two are used to determine the strength of decoupled learning of the model, and the latter is used to determine the strength of causal intervention.
@@ 5 Experiments and Analysis
5.1 Experimental Settings
5.1.1 Datasets
We evaluate our model and baselines on six datasets, including ICEWS14 [10], ICEWS18 [14], ICEWS05-15 [10], YAGO [24], WIKI [17] and GDELT [18]. Integrated Crisis Early Warning System (ICEWS) [3] contains international event information and is a commonly used benchmark dataset for TKGs link prediction. We choose three subsets of it, i.e., ICEWS14, ICEWS18, and ICEWS05-15, which contain events occurring in 2014, 2018, and 2005 to 2015 respectively. The WIKI and YAGO datasets are subsets of the Wikipedia history and YAGO3, respectively. The GDELT is from the Global data on events, location, and tone. To ensure a fair comparison, we use the split manner provided by Sun et al. [28]. Statistics of the datasets are summarized in Table 1.
@@ 5.1.1 Datasets
Baselines. For the link prediction task, we compare CEGRL-TKGR model with two categories of KGR models: (1) static KGR models, including TransE [2], DistMult [33], ComplEx [30] and R-GCN[26]. We apply these models in static KGs that ignore timestamp information. (2) TKGR models, including TTransE [17], TA-DistMult [10], TNTComplEx [16], RE-GCN [21], GHT [28], EvoKG [25], TITer[29], xERTE [12], TLogic[23] and CEN[20].
Evaluation Metrics. The mean reciprocal rank (MRR) and Hits@k are standard metrics for the TKG link prediction task. MRR is the average reciprocal of the correct query answer rank. Hits@$k$ indicates the proportion of correct answers among the top $k$ candidates. We used a more reasonable time-aware filter setting to report our results. The time-aware filtering setting filters out only the four groups that occur at query time, and can simulate extrapolated prediction tasks in the real world [29].
Implementation Details. We set the dimension of all embeddings and hidden states to 200. The number of layers of the R-GCN is set to 1 for YAGO and 2 for the other datasets. The optimal number of historical snapshots is set to 8, 10, 10, 1, 2, and 6 for ICEWS14, ICEWS18, ICEWS05-15, YAGO, WIKI, and GDELT, respectively. To fair comparison, static graph constraints are added for ICEWS14, ICEWS18, and ICEWS05-15. The channel number for decoding is set to 50, and the kernel size is set to 4$\times$3. We tried several different values for $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$, and finally chose 0.5, 0.5, 0.3. We use Adam to optimize the parameters, with a learning rate of 0.001. All of the experiments are processed on a Linux server with CPU Xeon Gold 6142, RAM 64G, and Nvidia 4090 GPU.
5.2 Experimental Results and Discussion
Table 2 and Table 3 report the experimental results of the link prediction task on six TKG datasets. Static KG embedding methods fell far behind CEGRL-TKGR due to their inability to capture temporal dynamics. Our method is also superior to other temporal knowledge graph extrapolation methods in predicting events. The improved performance shows that surface patterns and noise are widely present in several real-world data sets. The previous methods are generally inadequate in design. The model based on evolutionary representation will learn the inherent confounding features in TKG when gathering neighborhood information and transmitting historical information, and the model based on rule-based inference will mine the false correlation in the data, all of which will lead to the model making non-causal predictions in the inference stage. Our model incorporates causal theory into the TKG inference task and visibly separates causal features from confounding features. This helps to protect the model from surface patterns and noise present in the data and to uncover the real reasons that affect the formation of links between entities. TiTer and EvoKG show excellent performance on YAGO datasets because the former’s historical fact search strategy works well on smaller datasets, while the latter’s modeling of event timing works well on datasets containing events at relatively regular time intervals.
@@ 5.2 Experimental Results and Discussion
5.3 Performance on Noisy Temporal Knowledge Graphs
To explore whether CEGRL-TKGR can alleviate the problem of the TKGR model affected by noise and surface patterns in training data. Taking YAGO and WIKI datasets as examples, we respectively test the performance of our model and the model without the causal enhancement module under different noise deviations. Noisy TKGs were generated by randomly replacing a certain percentage of positive triples in the training set. The experimental results are shown in Fig. 4.
(a) MRR results on YAGO.
(b) Hits@1 results on YAGO.
(c) MRR results on WIKI.
(d) Hits@1 results on WIKI.
Figure 4: The performance of the CEGRL-TKGR model on the noisy dataset.
[…]
5.4 Ablation Study and Analysis on Parameter Sensitivity
@@
Ablation Study. In this subsection, we investigate the effectiveness of causally enhanced and time-interval guided decoders for link prediction. Specifically, CEGRL-TKGR w/o TD means that no time interval vector is used to guide the decoder to work, and CEGRL-TKGR w/o CE means that the model removes causal decoupling and causal intervention parts. Table 4 shows the results of ablation experiments, which indicate the effectiveness of these two components. As can be seen from the results in the table, for data sets such as YAGO and WIKI that contain relatively regular time intervals, a temporal gap-guided decoder can capture this time interval pattern well enough to make accurate predictions. At the same time, it does not degrade performance even for time-interval insensitive data sets. Our causal enhancement section, under the independent constraint of emphasizing causal features and confounding features, eliminates the influence of the fast bridge through causal intervention, forcing the model to learn the intrinsic causes of the events. It is worth noting that our causal enhancement module can be seen as a flexible component that can be easily used in many GNN-based TKGR frameworks.
@@
Table 4: Ablation study of CEGRL-TKGR model on the six datasets. ”w/o” means ”without”.
@@
Sensitivity of hyperparameters. In the proposed model, $\lambda_{1}$ and $\lambda_{2}$ jointly affect the disentanglement intensity of causal and confounding features, and $\lambda_{3}$ controls the intensity of causal intervention. We studied the sensitivity of parameters in different data sets. Specifically, one parameter is fixed at 0.5 and the other parameter varies in [0,1] with a step size of 0.1. As shown in Fig. 5, the model is relatively stable in most parameter selection cases, but on noisy data sets, the model has higher requirements for hyperparameters, and extreme values will degrade the performance of the model. The best range for $\lambda_{1}$, $\lambda_{2}$ is about 0.5 to 0.7. $\lambda_{3}$ should be a relatively small value, ranging from 0.3 to 0.6.
Figure 5: Sensitivity analysis of $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$ parameters.
5.5 Case Study
To show the ability of the CEGRL-TKGR model to learn the causal relationships that lead to events, we present a real case of CEGRL-TKGR versus a model that does not use causal augments in the ICEWS14 test set on Greek politics.
Figure 6: A Case study of the proposed CEGRL-TKGR model on the ICEWS14 test set.
As shown in Fig. 6, for inquiries (Movement Party Chairman, consult, ?). The model that does not use causal augmentation does learn the information in history, it learns the pattern of events that occurred multiple times in the training set, that is, at multiple different times (movement party chairman, consult, prime Minister), and gives the answer accordingly as the prime Minister. In contrast, CEGRL-TKGR was keen to capture events close to the inquiry (Movement Party chairman, express intent to meet, Alliance Party Chairman), which directly led to the occurrence of this event (Movement Party Chairman, consult, Alliance Party Chairman). The model is guided to make correct predictions by the learning of causality.
6 Conclusion
In this paper, we revisit the GNN-based TKGR model from the perspective of causality, on this basis, we propose an innovative causal enhanced graph representation learning framework. By synergistically integrating causal structures with graph representation learning in TKGs, CEGRL-TKGR overcomes the problem of existing models unintentionally learning biased data representations and mining for false correlations. Comprehensive experiments have proved the effectiveness of CEGRL-TKGR. In future work, we desire to explore how causal learning can help TKGR tasks in addition to representation learning.

### v2

@@ 4.3 Disentangled Causal and Confounding Features
In the previous subsection, the entity and relation representations are learned based on GNN-contained causal and confounding factors, and we separate them at the presentation level, which provides a solution to the previously mentioned problem of not being able to separate these two features at the data level. To do this, we introduce a decoupling module to decouple causal and confounding features. Taking the entity embedding matrix as an example, it is represented as follows:
@@
$\textbf{E}_{C}=\textbf{E}\odot\textbf{D}_{E,C},\textbf{E}_{N}=\textbf{E}\odot\textbf{D}_{E,N}.$ (7)
We want the two embeddings learned from the decoupling module to be as independent as possible, which is essential to accurately separate causal and confounding features Chen et al. (2023). Mutual information is a basic quantity to measure the nonlinear correlation of two random variables. Minimizing mutual information is a feasible scheme to decouple causal features from confounding features. Specifically, we implement this process with contrastive log-ratio upper-bound MI estimator Cheng et al. (2020); Wu et al. (2021), which utilizes variational distributions $q$ and a neural network to approximate the true distribution. We define the objective function as follows:
$\begin{split}\mathcal{L}_{mi}=&\ \mathbb{E}_{p(\textbf{E}_{C},\textbf{E}_{N})}\left[\log q_{\theta}(\textbf{E}_{N}|\textbf{E}_{C})\right]\\ &-\mathbb{E}_{p(\textbf{E}_{C})}\mathbb{E}_{p(\textbf{E}_{N})}\left[\log q_{\theta}(\textbf{E}_{N}|\textbf{E}^{\prime}_{C})\right].\end{split}$ (8)
We perform the same operation with relation embedding decoupling, after which we obtain $\textbf{R}_{C}$ and $\textbf{R}_{N}$.
@@ 4.4 Temporal Gap Guided Decoder
After the causal and confounding embeddings of entities and relations in the derived data, we use a specially crafted decoder to determine the likelihood score of potential entities and relations. Events or facts in a data stream may span different periods. For example, major political events may occur in rapid succession over a short period, while certain rare natural phenomena may occur sporadically and at longer intervals. With this in mind, it is reasonable to consider the time intervals of events to get an accurate picture of their temporal relationship. The key to the design of our decoder is the time interval vector, which guides the decoding process in considering the event time interval while calculating the fraction. Formulaic as:
$\mathbf{t}_{s}=\boldsymbol{\alpha}_{s}t+\boldsymbol{\beta}_{s},\ \mathbf{t}_{l}=\boldsymbol{\alpha}_{l}t+\boldsymbol{\beta}_{l}.$ (9)
Here, $\boldsymbol{\alpha}_{s},\boldsymbol{\beta}_{s},\boldsymbol{\alpha}_{l}$, and $\boldsymbol{\beta}_{l}$ signify learnable parameters. Adopting ConvTransE as our decoder, we introduce four variables, which traverse a one-dimensional convolutional layer followed by a fully connected layer, culminating in the extraction of a probability vector encompassing all entities. This process is mathematically articulated as:
$\begin{split}\mathbf{p}_{C}\left(e_{o}\mid e_{s},r,t\right)=&\ \operatorname{ReLU}\Big{(}\operatorname{ConvTransE}\big{(}\mathbf{e}_{s,C,t},\\ &\mathbf{r}_{C,t},\mathbf{t}_{s},\mathbf{t}_{l}\big{)}\Big{)}\mathbf{E}_{C,t}.\end{split}$ (10)
@@ 4.5 Causal Intervention and Training Objective
Causal-based embedding learns the intrinsic causes that cause events to occur, so the reasoning results obtained from causal-based embedding are expected to yield reasonable input results. We define the supervised classification loss as follows:
@@
$\begin{split}\mathcal{L}_{E,N}=&\ \frac{1}{\lvert\textbf{E}_{N,t}\rvert}\sum_{(e_{s},r,e_{o},t)\in\mathcal{G}}\operatorname{KL}\Big{(}\textbf{y}_{u},\\ &\ \log\mathbf{p}_{N}\left(e_{o}\mid e_{s},r,t\right)\Big{)},\end{split}$ (12)
@@
$\mathbf{p}_{I}\left(e_{o}\mid e_{s},r,t\right)$ $=\operatorname{ReLU}\Big{(}\operatorname{ConvTransE}\big{(}\mathbf{e}_{s,I,t},\mathbf{r}_{I,t},$
$\qquad\mathbf{t}_{s},\mathbf{t}_{l}\big{)}\Big{)}\mathbf{E}_{I,t},$ (14)
@@
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are designated as hyper-parameters, and the first two are used to determine the strength of decoupled learning of the model and the latter is used to determine the strength of causal intervention of the model.
@@ 5 Experiments and Analysis
5.1 Experimental Settings and Implementation Details
Datasets. We evaluate our model and baselines on six benchmark datasets, including ICEWS14 Garcia-Duran et al. (2018), ICEWS18 Jin et al. (2019), ICEWS05-15 Garcia-Duran et al. (2018), YAGO Mahdisoltani et al. (2014), WIKI Leblay and Chekol (2018) and GDELT Leetaru and Schrodt (2013). Statistics of the datasets are summarized in Table 1.
@@ 5.1 Experimental Settings and Implementation Details
Baselines. For the link prediction task, we compare CEGRL-TKGR model with two categories of KGR models: (1) static KGR models, including TransE Bordes et al. (2013), DistMult Yang et al. (2015), ComplEx Trouillon et al. (2016) and R-GCNSchlichtkrull et al. (2018). We apply these models in static KGs that ignore timestamp information. (2) TKGR models, including TTransE Leblay and Chekol (2018), TA-DistMult Garcia-Duran et al. (2018), TNTComplEx Lacroix et al. (2019), RE-GCN Li et al. (2021), GHT Sun et al. (2022), EvoKG Park et al. (2022), TITerSun et al. (2021), xERTE Han et al. (2021), TLogicLiu et al. (2022) and CENLi et al. (2022b).
Evaluation Metrics. The mean reciprocal rank (MRR) and Hits@k are standard metrics for the TKG link prediction task. MRR is the average reciprocal of the correct query answer rank. Hits@$k$ indicates the proportion of correct answers among the top $k$ candidates. We use a more reasonable time-aware filter setting to report our experimental results [^2] The time-aware filtering setting filters out only the four groups that occur at query time and can simulate extrapolated prediction tasks in the real world Sun et al. (2021)..
Implementation Details. The whole of training hyper-parameters and model configurations are summarized in Appendix A.1.
5.2 Experimental Results and Discussion
Table 2 and Table 3 report the experimental results of the link prediction task on six benchmark datasets. Static KG embedding methods fell far behind CEGRL-TKGR due to their inability to capture temporal dynamics in the TKG. Our method is also superior to other TKGR models in predicting events. The improved performance shows that surface patterns and noise are widely present in several real-world TKG datasets. The previous models are generally inadequate in design. CEGRL-TKGR based on evolutionary representation will learn the inherent confounding features in the TKG when gathering neighborhood information and transmitting historical information, and the model based on rule-based inference will mine the false correlation in the data, all of which will lead to the model-making non-causal predictions in the reasoning stage. Our model incorporates causal theory into the TKGR task and visibly separates causal features from confounding features. This helps to protect the model from surface patterns and noise present in the dataset and to uncover the real reasons that affect the formation of links between entities. TiTer and EvoKG show excellent performance on YAGO datasets because the former’s historical fact search strategy works well on smaller datasets, while the latter’s modeling of event timing works well on datasets containing events at relatively regular time intervals. More model configurations and experimental results are summarized in the Appendix.
6 Conclusion
In this paper, we revisit the GNN-based TKGR model from the causality perspective, on this basis, we propose a novel CEGRL-TKGR framework. By synergistically integrating causal structures with graph representation learning of the TKG, our framework overcomes the problem of existing TKGR models’ learning biased data representations and mining for false correlations unintentionally. Comprehensive experimental results and analysis have proved the effectiveness of CEGRL-TKGR.
Limitations and Future Work. The proposed CEGRL-TKGR is an innovational causal enhanced graph representation learning framework for optimizing feature representations directly using causal technology for the TGKR task. The limitations of CEGRL-TKGR are as follows:
  • From the dataset’s perspective, our research primarily focuses on TKG datasets, which may not verify the generalization ability of the CEGRL-TKGR framework to those time-interval insensitive graph datasets. Additionally, we aim to further conduct case studies to enhance the interpretability of the framework in the reasoning procedure as future work.
  • From the model’s perspective, our research evaluates the TKGR task alone. Theoretically, the GNN-based reasoning paradigm incorporated in the causal structure can be applied to any other graph representation learning tasks, such as triple classification Jaradeh et al. (2021), triple set prediction Zhang et al. (2024), and graph classification Liu et al. (2023). In future work, we desire to explore powerful disentanglement methods and more advanced causal intervention strategies to improve the CEGRL-TKGR’s performance for more rich graph representation learning-based tasks. Besides, the increased complexity of causal reasoning in the TKG is untouched.
  • From the adaptation’s perspective, to adapt the CEGRL-TKGR framework to more models, there are some hyper-parameters to control causal intervention and training. These hyper-parameters are sensitive to different models and datasets, hence it needs to take sufficient time to experiment to find the optimal values and combinations among them. Therefore, how to reduce the consumption in the above adaptation procedure upon the framework is worthy of consideration.
Acknowledgements. This work was supported by the National Natural Science Foundation of China (Grant No. 62202075, No. 62171111, No. 62376043), the Natural Science Foundation of Chongqing, China (No. CSTB2022NSCQ-MSX1404, No. CSTB2023NSCQ-MSX0091), Fundamental Research Funds for the Central Universities (No. SWU-KR24008), Science Research Project, Chongqing College of International Business and Economics (No. KYZK2024001), Key Laboratory of Data Science and Smart Education, Hainan Normal University, Ministry of Education (No. DSIE202206), and the State Key Laboratory of Public Big Data, Guizhou University (No. PBD2024-0501), Ming Liu’s research was funded in part by Australian Research Council Linkage (LP220200746). Yongbin Qin’s work was supported by the National Natural Science Foundation of China (No. 62166007, No. 62066008) and the National Key R&D Program of China (No. 2023YFC3304500).
Appendix A Appendix
A.1 Implementation Details
We set the dimension of all embeddings and hidden states to 200. The number of layers of the R-GCN is set to 1 for YAGO and 2 for the other datasets. The optimal number of historical snapshots is set to 8, 10, 10, 1, 2, and 6 for ICEWS14, ICEWS18, ICEWS05-15, YAGO, WIKI, and GDELT, respectively. To fair comparison, static graph constraints are added for ICEWS14, ICEWS18, and ICEWS05-15. The channel number for decoding is set to 50, and the kernel size is set to 4$\times$3. We try several different values for $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$, and finally chose 0.5, 0.5, 0.3. We use Adam to optimize the parameters, with a learning rate of 0.001. All of the experiments are processed on a Linux server with CPU Xeon Gold 6142, RAM 64G, and Nvidia 4090 GPU.
A.2 Ablation Study
@@
We investigate the effectiveness of causally enhanced and time-interval guided decoders for the link prediction task. Specifically, CEGRL-TKGR w/o TD means that no time interval vector is used to guide the decoder to work, and CEGRL-TKGR w/o CE means that the model removes causal decoupling and causal intervention parts. Table 4 shows the results of ablation experiments, which indicate the effectiveness of these two components. As can be seen from the results in the table, for datasets such as YAGO and WIKI that contain relatively regular time intervals, a temporal gap-guided decoder can capture this time interval pattern well enough to make accurate predictions. At the same time, it does not degrade performance even for time-interval insensitive datasets. Our causal enhancement module, under the independent constraint of emphasizing causal features and confounding features, eliminates the influence of the fast bridge through causal intervention, forcing the model to learn the intrinsic causes of the events. It is worth noting that our causal enhancement module can be seen as a flexible component that can be easily used in several GNN-based reasoning frameworks.
@@
Table 4: The ablation study of our model on the six benchmark datasets. "w/o" means "without".
@@
A.3 Parameter Sensitivity Analysis
In the CEGRL-TKGR, $\lambda_{1}$ and $\lambda_{2}$ jointly affect the disentanglement intensity of causal and confounding features, and $\lambda_{3}$ controls the intensity of causal intervention. We study the sensitivity of parameters in different benchmark datasets, as depicted in Fig. LABEL:fig:hyper. Specifically, one parameter is fixed at 0.5 and the other parameter varies in [0,1] with a step size of 0.1. The model is relatively stable in most parameter selection cases, but on noisy datasets, the model has higher requirements for hyper-parameters, and extreme values will degrade the performance of the model. The best range for $\lambda_{1}$, $\lambda_{2}$ is about 0.5 to 0.7. $\lambda_{3}$ should be a relatively small value, ranging from 0.3 to 0.6.
A.4 Performance on Noisy Temporal Knowledge Graphs
To explore whether the proposed CEGRL-TKGR can alleviate noise and surface patterns, we randomly replace a certain percentage of positive triples in the training set of each TKG dataset in form of noisy TKGs. Taking YAGO and WIKI datasets as examples, we test the performance of CEGRL-TKGR and CEGRL-TKGR w/o CE under different noise deviations, respectively. The experimental results are shown in Fig. LABEL:fig:noise.
From the experimental results, we can draw the following conclusion: when the noise in the dataset increases, the performance of models lacking the recognition of causal features and confounding features will deteriorate sharply, and the performance of MRR and Hits@1 will decrease, which indicates that the CEGRL-TKGR w/o CE is easy to capture data bias and make wrong predictions based on it. In contrast, CEGRL-TKGR uses the causal enhancement module to effectively reduce the impact of confounding features and shows more stable performance on these two noisy TKG datasets. The performance degradations on MRR and Hits@1 are significantly smaller than those without the causal module.